Three objects are located along the x-axis. Object A with a charge of +5.6 µC is located at the origin. Object B has a charge of -4.2 µC and is located at the -1.2 m location. Object C has a charge of +7.7 µC and is located at the +2.4 m location. Determine the magnitude and direction of the net electric force acting upon object A.



1
Expert's answer
2022-01-10T09:10:25-0500

The force is


"F=F_{AB}+F_{AC},\\\space\\
F=\dfrac1{4\pi \epsilon_0}\bigg(\dfrac{q_Aq_B}{r_{AB}^2}+\dfrac{q_Aq_C}{r_{AC}^2}\bigg)=0.21\text{ N}."